# Sensitive 3mm Imaging of Discrete Sources in the Fields of tSZ-Selected Galaxy Clusters

### v1

@@ Abstract
In this paper we present the results of a blind survey for compact sources in 243 Galaxy clusters that were identified using the thermal Sunyaev-Zeldovich effect (tSZ). The survey was carried out at 90 GHz using MUSTANG2 on the Green Bank telescope and achieved a $5\sigma$ detection limit of 1 mJy in the center of each cluster. We detected 24 discrete sources. The majority (18) of these correspond to known radio sources, and of these, 5 show signs of significant variability, either with time or in spectral index. The remaining sources have no clear counterparts at other wavelengths. Searches for galaxy clusters via the tSZ effect strongly rely on observations at 90 GHz, and the sources found have the potential to bias mass estimates of clusters. We compare our results to the simulation Websky that can be used to estimate the source contamination in galaxy cluster catalogs. While the simulation showed a good match to our observations at the clusters’ centers, it does not match our source distribution further out. Sources over ${}^{\prime\prime}$ from a cluster’s center bias the tSZ signal high, for some of our sources, by over 50%. When averaged over the whole cluster population the effect is smaller but still at a level of 1 to 2 percent. We also discovered that unlike previous measurements and simulations we see an enhancement of source counts in the outer regions of the clusters and fewer sources than expected in the centers of this tSZ selected sample.
@@
facilities: Green Bank Telescope (MUSTANG2), WISE, SPIREsoftware: astropy (Astropy Collaboration et al., 2013, 2018)
@@ 1 Introduction
Galaxy clusters are the most massive virialized structures in the universe, and they constitute extremely useful laboratories for probing cosmology. For example, cluster counts have been used to estimate cosmological parameters (e.g., Komatsu et al., 2011) and can help discern between different models for the nature of dark energy (Allen et al., 2011). Optical and X-ray surveys have long been used to detect and study galaxy clusters. More recently, millimeter-wave surveys have been used to find clusters by the thermal Sunyaev-Zeldovich effect (tSZ) (see Mroczkowski et al., 2019, for a review). The tSZ is a spectral distortion of the Cosmic Microwave Background (CMB) due to inverse Compton scattering off hot electrons in the intracluster medium (ICM). As this is nearly redshift-independent, the tSZ is a powerful tool for detecting galaxy clusters. Experiments such as the Atacama Cosmology Telescope (ACT; Thornton et al., 2016) and the South Pole Telescope (SPT; McMahon et al., 2009) have discovered many clusters. The latest catalogs from the SPT contains over 600 optically confirmed clusters (Bleem et al., 2020, 2023) while Data Release 5 (DR5) from ACT contains over 4000 (Hilton et al., 2021). In the future, the Simons Observatory and CMB-S4 are expected to discover an order of magnitude more clusters (Ade et al., 2019; Abazajian et al., 2019).
@@
In order to be used as cosmological tools, it is necessary to know each cluster’s mass. The magnitude of the tSZ effect, the Compton-$y$ parameter, is proportional to the integrated line-of-sight pressure and its sum over the cluster, Y, is a good proxy for the cluster mass. For large surveys the sum over the cluster is typically calculated using multi-frequency matched filters normalized such that the peak at the cluster location is equal to Y. This is the centralized Compton-$y$ parameter and in the ACT DR5 cluster catalog is refereed to as $y_{c}$ (Hilton et al., 2021). There is however, an 11% scatter in the ACT Compton-$y$ to mass relation (Y-M)(Battaglia et al., 2012). For tSZ galaxy cluster surveys to fulfill their promise as precision cosmology tools, it is crucial to understand these systematic effects. The mass selection effects in tSZ galaxy cluster surveys are well understood and simulations have been used to predict the effects of cluster mergers (Yu et al., 2015). However, past studies disagree on the effect that compact radio sources have on the detection of galaxy clusters and how much this might contribute to the scatter in Y-M relationships (Coble et al., 2007; Gralla & Marriage, 2020).
@@
Although the mass of clusters can be found by weak lensing, the weak lensing signal is not detectable in a single cluster. It is necessary to stack clusters with similar masses (or measured Y values). Scatter in measured Y has the potential to bias weak lensing measurements as some clusters in each bin could be of significantly different mass.
@@
Dicker et al. (2021, herafter D2021) used an ad-hoc sample of galaxy clusters observed at 90 GHz by MUSTANG2 on the Green bank telescope to investigate the effects of compact sources on the measured $y_{c}$. This sample of 30 clusters had been selected for deep observations by different projects for a range scientific aims. Approximately one third of the clusters had compact sources brighter than 1 mJy, enough to bias mass estimates by 2–100%. However, these clusters were selected for study due to their interesting properties (e.g. known mergers) which could bias the results. This motivated the study in this paper of a search for sources in a more representative sample of clusters.
@@ 1.1 Compact sources
To maximize mapping speed, tSZ survey instruments have a beam size comparable to the angular size of clusters ($1.4^{\prime}$$2.2^{\prime}$ for ACT). At these resolutions compact sources can blend in with the cluster. The sources can be contained within the cluster or be in the foreground or background. Most of the sensitivity of tSZ measurements comes from a combination of data taken in the 90 GHz and 150 GHz atmospheric windows where the tSZ signal is negative. As shown in Figure 1, a source with a projected location near the cluster center cancels out some of this signal resulting in a decrease in measured cluster mass. When a source has a line of sight location further from the cluster center, the matched filters in the data pipelines used to find clusters can result in an increase in the measured mass of a cluster (D2021). The magnitude of the effect depends on the flux of the source at each tSZ band (90 and 150 GHz) which can also be expressed as a flux at 90 GHz and a spectral index $\alpha$ between the two. These spectral indexes can vary widely depending on source type and are often different from spectral indexes calculated at lower frequencies. Following the notation of D2021, the effect of a source with a 90 GHz flux $I$ on the measured centralized ACT Compton-$y$ ($y_{c}$) is given by:
@@
where $\delta_{\tilde{y}_{0}}=-8.76{\times}10^{-6}$ is a constant equal to the effect on the Compton-$y$ of a 1 mJy source at 90 GHz with a spectral index of $-0.7$.
Figure 1: How a source of fixed amplitude at 90 GHz affects the recovered Compton-$y$ parameter varies depending on its projected distance from the cluster center and the source spectral index. [Reproduced from Dicker et al. (2021)]. This can be represented by the functions $N(r)$ (upper panel) and $A(\alpha)$ (lower panel). The exact shape depends on the matched filters used by a given tSZ survey but will be similar to these examples from ACT DR5.
This relationship, found by passing clusters with simulated sources through the real ACT DR5 cluster pipeline  [^1] https://nemo-sz.readthedocs.io/en/latest/, can be used to marginalize over the effects of point sources - but only if the distribution in flux in the tSZ bands, the distribution in spectral index and typical locations within a cluster are known. Much of our information on compact sources in clusters comes from low frequency (1–4 GHz) radio surveys. To predict the effect of point sources on cluster mass determination would therefore require flux extrapolations over 1-2 orders of magnitude. In addition, near to 90 GHz, it is common for radio sources to have breaks in their spectral indexes (Tucci et al., 2011) and emission from dust can start to become important. The principle reason for uncertainty in the effects of compact sources on the Y-M relationship is due to the limited number of direct observations at tSZ frequencies. In this paper we present results from a large survey for compact sources in tSZ selected clusters. Our observations and data reduction are in Sections 2 and 3. Section 4 describes our search for counterparts to the sources we found at other wavelengths and in Section 5 we compare our results with the Websky model (Stein et al., 2020) for point source distributions and the smaller cluster sample in D2021. Our conclusions can be found in Section 6.
@@ 2 Observations
We surveyed all clusters in the ACT DR5 sample above a declination of $-6^{\circ}$ and within Right Ascensions between $8^{h}40^{m}$$13^{h}00^{m}$ and $2^{h}00^{m}$$3^{h}00^{m}$ at $9^{\prime\prime}$ angular resolution in the continuum at 90 GHz. These targets have good visibility from the Green Bank Telescope (GBT) at night during the high frequency observing season. To aid with the classification of sources, there is good coverage from other surveys at higher and lower frequencies including: WISE (Schulz et al., 2018a), the Faint Images of the Radio Sky at Twenty-cm (FIRST) point source catalog (White et al., 1997), and the VLA Sky Survey (VLASS) (Cohen et al., 2007).
@@
Within the survey footprint there are 300 ACT-identified clusters. No other cuts were applied to the sample so our survey accurately reflects the population of ACT DR5 tSZ-selected galaxy clusters. One notable aspect of the ACT DR5 cluster sample is that it excludes any targets on sight-lines to bright point sources ($\geq 10\,{\rm mJy}$ at 150 GHz). We refer to this survey as the M2-ACT cluster survey.
@@
It is important that the survey be made at angular resolutions better than ${}^{\prime\prime}$ so that sources can be separated from the typical core size of a galaxy cluster. Sources up to $4^{\prime}$ from cluster centers can still have an effect on the measured tSZ (see Figure 1), consequently an ideal survey to quantify the effects of point sources would measure a large number of clusters out to this radius, significantly larger than the field-of-view of ALMA. The MUSTANG2 camera (M2) (Dicker et al., 2014) on the Green Bank Telescope has $4.2^{\prime}$ field-of-view, a resolution of $9^{\prime\prime}$ and can reach noise levels below 60 $\mu$Jy/beam in an hour. MUSTANG2 also has a bandpass well matched to CMB experiments such as ACT thus simplify data analysis making it an ideal instrument for this survey.
@@
The M2-ACT cluster survey target depth was a compromise between the desire to detect faint sources which, while common, only affect the measured mass of a cluster by a few percent and observing enough clusters in our allotted telescope time to gain reliable statistics on less common brighter sources which affect the measured mass by far more. Scaling from the deep observations in Dicker et al. (2021), in 300 clusters we could expect 90 sources brighter than 1 mJy. A 1 mJy source with a line-of-sight location in the center of a cluster would give a 12% change in measured Compton-$y$ for a typical ($y_{c}=7.2\times 10^{-5}$) ACT DR5 cluster. Consequentially our target noise was set to be at 200 $\mu$Jy to enable $5\sigma$ detections of 1 mJy sources in the center of each ACT DR5 cluster. For MUSTANG2 this depth can be achieved with just 5 minutes per cluster. Deeper observations would mean significantly fewer clusters could be included in the survey. A wider, shallower survey was not possible due to limits of the GBT slew rate and the desire to observe each region of the cluster multiple times with independent detectors. A scan pattern was selected which, in combination with the MUSTANG2 camera’s field of view, would provide useful coverage out to ${}^{\prime}$ radius for each target. This comfortably encompasses the region within which discrete sources can affect the ACT DR5 cluster selection.
@@
Table 1: A partial list of ACT DR5 clusters with successful observations and the depth of each map. Compton-$y$ values were taken to be the $y_{c}$ values in the DR5 catalog. The full table can be found as an on-line extra.
@@ 3 Data reduction
The standard MIDAS data reduction pipeline (see Romero et al., 2020, for details) was used to produce maps for each of our targets. Simulations described below showed that moderately aggressive high pass filtering of the raw MUSTANG2 timestreams at 0.05 Hz produced the best recovery of point sources. Extra care had to be taken when flagging detector glitches to avoid false positives. Inclement weather resulted in higher than anticipated noise in some clusters but 243 clusters had noise levels better than our 0.2 mJy RMS target (Table 1). For automatic detection of compact sources a matched filter to the MUSTANG2 beam would be ideal. However, as the noise properties in the maps varies (both between maps and as a function of radius), such a filter is challenging to calculate. Instead we used a Difference of Gaussians (DoG) filter which approximates a matched filter by convolving with two different width Gaussians ($\sigma_{1}=3.6^{\prime\prime}$ and $\sigma_{2}=13.6^{\prime\prime}$) and taking the difference. These widths were found empirically to produce the strongest detections of point sources injected into our maps. As well as a regular map, a noise map was made by flipping the sign of half of the data from each cluster. Using the noise map normalized by the integration time at each point, it was possible to make signal-to-noise (SNR) maps for each cluster. Sources can then be found by looking for positive peaks. Once a source location is established, a fit for its amplitude was carried out in the original signal map using a 2D Gaussian.
Figure 2: Histograms of the signal-to-noise ratio in the MUSTANG2 maps. Inside a radius of $3^{\prime}$ the noise is Gaussian and a $5\sigma$ cut should produce less than one false positive in our survey. Outside this radius the coverage of the maps was lower and the noise statistics non-Gaussian, consequently a $7\sigma$ cut was needed to obtain the same false positive rate. The excess power on the positive side is due to sources and is not present if these regions are excluded.
@@
Initial searches for sources used a uniform $5\sigma$ cut. However, outside of a radius of $3^{\prime}$ the number of sources increased rapidly (even though these areas of each map had higher noise). Most of these sources had a peak SNR below $7\sigma$. Closer examination showed the calculated noise in these areas of the map was, due to the lower number of observations per pixel, non-Gaussian (Figure 2). However, sources at $r>3^{\prime}$ have five times less impact on the measured mass of the cluster (Figure 1) so accepting a slightly higher threshold at $r>3^{\prime}$ of $7\sigma$ is still consistent with our goal of of characterizing the population all sources that could significantly affect the measured Compton-$y$ of a typical cluster.

### v2

# Sensitive 3 mm Imaging of Discrete Sources in the Fields of Thermal Sunyaev–Zel’dovich Effect-Selected Galaxy Clusters

@@ Abstract
In this paper, we present the results of a blind survey for compact sources in 243 galaxy clusters that were identified using the thermal Sunyaev-Zel’dovich effect (tSZ). The survey was carried out at 90 GHz using MUSTANG2 on the Green Bank Telescope and achieved a $5\sigma$ detection limit of 1 mJy in the center of each cluster. We detected 24 discrete sources. The majority (18) of these correspond to known radio sources, and of these, five show signs of significant variability, either with time or in spectral index. The remaining sources have no clear counterparts at other wavelengths. Searches for galaxy clusters via the tSZ effect strongly rely on observations at 90 GHz, and the sources found have the potential to bias mass estimates of clusters. We compare our results to the Websky simulation that can be used to estimate the source contamination in galaxy cluster catalogs. While the simulation showed a good match to our observations at the clusters’ centers, it does not match our source distribution further out. Sources over 104′′ from a cluster’s center bias the tSZ signal high, for some of the sources found, by over 50%. When averaged over the whole cluster population, the effect is smaller but still at a level of 1 to 2%. We also discovered that unlike previous measurements and simulations, we see an enhancement of source counts in the outer regions of the clusters and fewer sources than expected in the centers of this tSZ-selected sample.
@@
facilities: GBT (MUSTANG2), WISE, Herschel (SPIRE)software: astropy (Astropy Collaboration et al., 2013, 2018)
@@ 1 Introduction
Galaxy clusters are the most massive virialized structures in the Universe, and they constitute extremely useful laboratories for probing cosmology. For example, cluster counts have been used to estimate cosmological parameters (e.g., Komatsu et al., 2011) and can help discern between different models for the nature of dark energy (Allen et al., 2011). Optical and X-ray surveys have long been used to detect and study galaxy clusters. More recently, millimeter-wave surveys have been used to find clusters by the thermal Sunyaev-Zel’dovich effect (tSZ, see Mroczkowski et al., 2019, for a review). The tSZ is a spectral distortion of the Cosmic Microwave background (CMB) due to inverse Compton scattering off hot electrons in the intracluster medium (ICM). As this is nearly redshift-independent, the tSZ is a powerful tool for detecting galaxy clusters. Experiments such as the Atacama Cosmology Telescope (ACT; Thornton et al., 2016) and the South Pole Telescope (SPT; McMahon et al., 2009) have discovered many clusters. The latest catalogs from the SPT contain over 600 optically confirmed clusters (Bleem et al., 2020, 2024), while Data Release 5 (DR5) from ACT contains over 4000 (Hilton et al., 2021). In the future, the Simons Observatory and CMB-S4 are expected to discover an order of magnitude more clusters (Ade et al., 2019; Abazajian et al., 2019).
1.1 Clusters and Cosmology
@@
In order to be used as cosmological tools, it is necessary to know each cluster’s mass. The magnitude of the tSZ effect, the Compton-$y$ parameter, is proportional to the integrated line-of-sight pressure, and its sum over the cluster, $Y$, is a good proxy for the cluster mass. For large surveys, the sum over the cluster is typically calculated using multifrequency matched filters normalized such that the peak at the cluster location is equal to $Y$. This is the centralized Compton-$y$ parameter and in the ACT DR5 cluster catalog is refereed to as $y_{c}$ (Hilton et al., 2021). There is however, an 11% scatter in the ACT Compton-$y$ to mass relation ($Y$-$M$; Battaglia et al., 2012). For tSZ galaxy cluster surveys to fulfill their promise as precision cosmology tools, it is crucial to understand these systematic effects. The mass selection effects in tSZ galaxy cluster surveys are well understood and simulations have been used to predict the effects of cluster mergers (Yu et al., 2015). However, past studies disagree on the effect that compact radio sources have on the detection of galaxy clusters and how much this might contribute to the scatter in $Y$-$M$ relationships (Coble et al., 2007; Gralla & Marriage, 2020).
@@
Although the mass of clusters can be found by weak lensing, the weak-lensing signal is not detectable in a single cluster. It is necessary to stack clusters with similar masses (or measured Y values). Scatter in measured Y has the potential to bias weak lensing measurements as some clusters in each bin could be of significantly different mass.
@@
Dicker et al. (2021, hereafter D2021) used an ad hoc sample of galaxy clusters observed at 90 GHz by MUSTANG2 on the Green Bank Telescope (GBT) to investigate the effects of compact ($<15^{\prime\prime}$ in diameter) sources on the measured $y_{c}$. This sample of 30 clusters had been selected for deep observations by different projects for a range of scientific aims. Approximately one third of the clusters had compact sources brighter than 1 mJy, enough to bias individual mass estimates by 2%–100%. However, these clusters were selected for study due to their interesting properties (e.g. known mergers), which could bias the results. This motivated the study in this paper of a search for sources in a more representative sample of clusters.
@@ 1.2 Compact sources
[…]
@@
where $\delta_{\tilde{y}_{0}}=-8.76{\times}10^{-6}$ is a constant equal to the effect on the Compton-$y$ of a 1 mJy source at 90 GHz, $r=0$, and with a spectral index of $-0.7$, and $N(r)$ and $A(\alpha)$ are functions representing the response to a source as a function of projected radius ($r$) from the cluster center and spectral index ($\alpha$).
Figure 1: How a source of fixed amplitude at 90 GHz affects the recovered Compton-$y$ parameter varies depending on its projected distance from the cluster center and the source spectral index. [Reproduced from Dicker et al. (2021)]. This can be represented by the functions $N(r)$ (upper panel) which has a null at 104′′ and $A(\alpha)$ (lower panel). The exact shape depends on the matched filters used by a given tSZ survey but will be similar to these examples from ACT DR5.
This relationship, found by passing clusters with simulated sources through the real ACT DR5 cluster pipeline, [^1] https://nemo-sz.readthedocs.io/en/latest/ can be used to marginalize over the effects of compact sources — but only if the distribution in 90 GHz flux, the distribution in spectral index and the typical locations within a cluster are known. Much of our information on compact sources in clusters comes from low-frequency (1–4 GHz) radio surveys. To predict the effect of compact sources on cluster mass determination would therefore require flux extrapolations over 1–2 orders of magnitude. In addition, near to 90 GHz, it is common for radio sources to have breaks in their spectral indexes (Tucci et al., 2011) and emission from dust can start to become important. The principal reason for uncertainty in the effects of compact sources on the $Y$-$M$ relationship is due to the limited number of direct observations at tSZ frequencies.
In this paper, we present results from a large survey for compact sources in tSZ-selected clusters. Our observations and data reduction are in Sections 2 and 3. Section 4 describes our search for counterparts to the sources we found at other wavelengths and in Section 5 we compare our results with the Websky model (Stein et al., 2020) for source distributions and the smaller cluster sample in D2021. How the $Y$-$M$ relationship for ACT DR5 is affected by compact sources is covered in Section 6 and our conclusions can be found in Section 7. We report all errors at a $1\sigma$ / 68.5% confidence level.
@@ 2 Observations
We surveyed all clusters in the ACT DR5 sample above a declination of $-6^{\circ}$ and within Right Ascensions between $8^{h}40^{m}$$13^{h}00^{m}$ and $2^{h}00^{m}$$3^{h}00^{m}$ at $9^{\prime\prime}$ angular resolution in the continuum at 90 GHz. These targets have good visibility from the GBT at night during the high-frequency observing season. To aid in the classification of sources, there is good coverage from other surveys at higher and lower frequencies including the Wide-field Infrared Survey Explorer (WISE; Cutri et al., 2012), the Faint Images of the Radio Sky at Twenty-cm (FIRST) point source catalog (White et al., 1997), and the VLA Sky Survey (VLASS; Gordon et al., 2021).
@@
Within the survey footprint there are 300 ACT-identified clusters. No other cuts were applied to the sample, so our survey accurately reflects the population of ACT DR5 tSZ-selected galaxy clusters. One notable aspect of the ACT DR5 cluster sample is that it excludes any targets on sight lines to bright point sources ($\geq 10\,{\rm mJy}$ at 150 GHz). We refer to this survey as the M2-ACT cluster survey.
@@
It is important that the survey be made at angular resolutions better than $\sim$15′′ so that sources can be separated from the larger core size of a galaxy cluster. Sources up to $4^{\prime}$ from cluster centers can still have an effect on the measured tSZ (see Figure 1), consequently, an ideal survey to quantify the effects of compact sources would measure a large number of clusters out to this radius, significantly larger than the field-of-view of ALMA. MUSTANG2 (M2) (Dicker et al., 2014), a camera on the GBT has a $4.^{\prime}2$ field-of-view, a resolution of $9^{\prime\prime}$ and can reach noise levels below 60 $\mu$Jy/beam in an hour. MUSTANG2 also has a bandpass well matched to CMB experiments such as ACT thus simplifying data analysis and making it an ideal instrument for this survey.
@@
The M2-ACT cluster survey target depth was a compromise between the desire to detect faint sources that, while common, only affect the measured mass of a cluster by a few percent and observing enough clusters in our allotted telescope time to gain reliable statistics on less common brighter sources that affect the measured mass by far more. To estimate the expected number of sources in the M2-ACT cluster survey, one can scale from the deeper observations of 30 clusters in D2021 which found nine sources brighter than 1 mJy. If the two cluster samples have similar source populations, $90\pm 38$ sources of this strength can expected in 300 clusters. The relatively large errors come from the small D2021 sample size and in this calculation a normal distribution is assumed. From Equation 1, a 1 mJy source with a line-of-sight location in the center of a cluster would give a 12% change in measured Compton-$y$ for a typical ($y_{c}=7.2\times 10^{-5}$) ACT DR5 cluster. Consequentially, our target noise was set to be at 200 $\mu$Jy to enable $5\sigma$ detections of 1 mJy sources in the center of each ACT DR5 cluster. For MUSTANG2, this depth can be achieved with just 5 minutes per cluster. Deeper observations would mean that significantly fewer clusters could be included in the survey. A wider, shallower survey was not possible due to the limits of the GBT slew rate and the desire to observe each region of the cluster multiple times with independent detectors. A scan pattern was selected that, in combination with the MUSTANG2 camera’s field-of-view, would provide useful coverage out to $\sim$4 in radius for each target. This comfortably encompasses the region within which discrete sources can affect the ACT DR5 cluster selection.
@@
Table 1: A partial list of ACT DR5 clusters with successful observations and the depth of each map. Compton-$y$ values were taken to be the $y_{c}$ values in the DR5 catalog. The full table can be found as supplementary material and individual maps can be found on https://doi.org/10.7910/DVN/FQNKYX (catalog dataverse).
@@ 3 Data reduction
The standard MUSTANG2 data reduction pipeline (MIDAS; see Romero et al., 2020, for details) was used to produce maps with 2′′ pixels for each of our targets. The simulations described below showed that moderately aggressive high-pass filtering of the raw MUSTANG2 time streams at 0.05 Hz produced the best recovery of compact sources. Extra care had to be taken when flagging detector glitches to avoid false positives. Inclement weather resulted in higher than anticipated noise in some clusters but 243 clusters had noise levels better than our 0.2 mJy RMS target (Table 1). For automatic detection of compact sources, a matched filter to the MUSTANG2 beam would be ideal. However, as the noise properties in the maps vary (both between maps and as a function of radius), such a filter is challenging to calculate. Instead, we used a difference-of-Gaussians (DoG) filter which approximates a matched filter by convolving with two different width Gaussians ($\sigma_{1}=3.^{\prime\prime}6$ and $\sigma_{2}=13.^{\prime\prime}6$) and taking the difference. These widths were found empirically to produce the strongest detections of point sources injected into our maps. As well as a regular map, a noise map was made by flipping the sign of half of the data from each cluster. Using the noise map normalized by the integration time at each point, it was possible to make signal-to-noise ratio (SNR) maps for each cluster. The signal from sources will always be positive, and for our target survey depth at MUSTANG2’s resolution, the negative signal from the galaxy clusters is negligible. This allows sources to be found by looking for positive peaks in the SNR maps. Once a source location is established, a fit for its amplitude and width was carried out in the original signal map using a 2D Gaussian. By using the original signal map for this fit, there was no need to account for the effects of the DoG filter.
Figure 2: Histograms of the signal-to-noise ratios of pixels in the MUSTANG2 maps. Inside a radius of $3^{\prime}$ the histograms fall off rapidly and a $5\sigma$ cut should produce less than one false positive in our survey which has $\mathcal{O}10^{5}$ independent beams. Outside this radius, there are fewer observations per pixel and the noise statistics have a significant non-Gaussian tail. Consequently, a $7\sigma$ cut was needed to obtain the same false-positive rate. The excess power on the positive side is due to compact sources and is not present if these regions are excluded.
@@
With any survey, the signal-to-noise cut is an important choice. A value too low will result in false detections. A value higher than the lowest cut needed to avoid false detections will result in valid sources being missed. The statistical properties of the pixels in the SNR maps are a good fit to a Gaussian with $\sigma=1$. The M2-ACT survey has approximately $10^{5}$ independent beams so, if this fit were perfect, a $5\sigma$ cut should give less than one false detection in the entire survey. Initial searches for sources used this uniform $5\sigma$ cut. However, outside of a radius of $3^{\prime}$ the number of sources found increased rapidly (even though these areas of each map had higher noise). Most of these sources had a peak SNR below $7\sigma$. Closer examination showed that, due to the lower number of observations in a few pixels, the calculated noise in these areas of the maps has a long non-Gaussian tail (Figure 2). From the negative side of the distribution in SNR values it can be seen that, between 3 and 4, a $7\sigma$ threshold is sufficient for the expected number of false detections to be less than one. Moreover, sources at $r>3^{\prime}$ have five times less impact on the measured mass of the cluster (Figure 1) so accepting a slightly higher threshold at $r>3^{\prime}$ of $7\sigma$ is still consistent with our goal of of characterizing the population all sources that could significantly affect the measured Compton-$y$ of a typical cluster.